OpenStudy (anonymous):

Find the point on the line (5,2,3)+t(2,3,1) closest to the point M(1,-1,-2)

OpenStudy (shadowfiend):

Hm. So you want to get these three equations as close as possible to true: 5 + 2t = 1 2 + 3t = -1 3 + t = -2 Basically, you want to minimize: 5 + 2t - 1 2 + 3t + 1 3 + t + 2 To be as close to 0 as possible. Does that help?

OpenStudy (anonymous):

Yes, thanks.